# A Reference Architecture for Governance of Cloud Native Applications

### v2

@@ VII Conclusion
This paper explores the realm of CNAs and the challenges involved in governing them. We present a reference architecture for robust CNA governance, applicable to both single and multi-cloud environments. Our architecture embeds governance directly into CNA design, offering a “batteries-included” approach that eases compliance for cloud professionals, allowing them to concentrate on product development rather than governance complexities. In addition, we illustrate the practical application of the architecture in different cloud setups, demonstrating its flexibility and ease of customization for specific governance needs. We seek to improve CNA management and operational efficiency in both regulated and non-regulated industries, offering a modular and adaptable governance framework valuable to both practitioners and researchers as a foundation for standardized CNA governance practices.
@@ C-F2 Leg Definition and Results
  • AWS Leg 1 $=$ RG_1_API_Gateway_timestamp -- cna_timestamp,
  • AWS Leg 2 $=$ RG_1_SQS_Forwarder_timestamp -- RG_1_API_Gateway_timestamp,
  • AWS Leg 3 $=$ RG_GOV_IMS_API_Gateway_timestamp -- RG_1_SQS_Forwarder_timestamp,
  • AWS Leg 4 $=$ RG_GOV_IMS_Converter_timestamp -- RG_GOV_IMS_API_Gateway_timestamp,
  • AWS Leg 5 $=$ RG_GOV_IMS_Archiver_timestamp -- RG_GOV_IMS_Converter_timestamp,
  • IBM Leg 1 $=$ RG_GOV_IMS_API_Gateway_timestamp -- cna_timestamp,
  • IBM Leg 4 $=$ RG_GOV_IMS_Converter_timestamp -- RG_GOV_IMS_API_Gateway_timestamp,
  • IBM Leg 5 $=$ RG_GOV_IMS_Archiver_timestamp -- RG_GOV_IMS_Converter_timestamp,

### v3

@@ VI Related Literature
Given the interdisciplinary nature of this work and the wide range of adjacent research domains, we have limited our literature review to peer-reviewed studies that focus specifically on CNA-related monitoring, observability, compliance, and governance. This targeted scope ensures relevance to the core components of our proposed architecture and avoids diluting the discussion with broader cloud or software engineering topics not directly aligned with CNA governance.
@@ VII Conclusion
This paper explores the realm of CNAs and the challenges involved in governing them. We present a reference architecture for robust CNA governance that is applicable across a range of single and multi-cloud environments. Rather than proposing a one-size-fits-all model, the architecture is designed to be modular and adaptable, enabling integration with diverse deployment contexts and enterprise governance policies. By embedding governance directly into CNA design, the framework offers a “batteries-included” approach that streamlines compliance efforts and enables cloud professionals to focus on core product development. The architecture’s real-world instantiation illustrates its flexibility and demonstrates how it may be tailored to meet specific organizational, regulatory, or operational needs. While this work highlights broad applicability, we also recognize that cloud environments vary in scale, tooling, and governance maturity. As such, additional customization may be required in some contexts. Overall, the proposed architecture aims to improve CNA management and operational efficiency, serving as a practical foundation for further research and standardization in cloud-native governance.
@@
Acknowledgments
This work was partially supported by the Natural Sciences and Engineering Research Council of Canada (Grant # RGPIN-2022-03886) and the IBM Centre for Advanced Studies (Project #1142).
@@ C-F2 Leg Definition and Results
  • AWS Leg 1 == RG_1_API_Gateway_timestamp - cna_timestamp,
  • AWS Leg 2 == RG_1_SQS_Forwarder_timestamp - RG_1_API_Gateway_timestamp,
  • AWS Leg 3 == RG_GOV_IMS_API_Gateway_timestamp - RG_1_SQS_Forwarder_timestamp,
  • AWS Leg 4 == RG_GOV_IMS_Converter_timestamp - RG_GOV_IMS_API_Gateway_timestamp,
  • AWS Leg 5 == RG_GOV_IMS_Archiver_timestamp - RG_GOV_IMS_Converter_timestamp,
  • IBM Leg 1 == RG_GOV_IMS_API_Gateway_timestamp - cna_timestamp,
  • IBM Leg 4 == RG_GOV_IMS_Converter_timestamp - RG_GOV_IMS_API_Gateway_timestamp,
  • IBM Leg 5 == RG_GOV_IMS_Archiver_timestamp - RG_GOV_IMS_Converter_timestamp,